# DL-Chain: Scalable and Stable Blockchain Sharding with High Concurrency via Dual-Layer Consensus

### v1

@@ Abstract
Sharding enhances blockchain scalability by partitioning nodes into multiple groups for concurrent transaction processing. Configuring a large number of small shards helps improve the transaction concurrency of a sharding system. However, it increases the fraction of malicious nodes within each shard, easily leading to shard corruption and jeopardizing system security. Some existing works have attempted to improve concurrency by reducing the shard size while maintaining security. However, they often require frequent and time-consuming recovery of corrupted shards, leading to severe system stagnation. Also, they usually require network-wide consensus to guarantee security, which limits scalability.
To address these issues, we propose DL-Chain, a blockchain sharding system that can securely provide high concurrency with stable and scalable performance. Our core idea is a Dual-Layer architecture and consensus, which consists of numerous smaller proposer shards (PSs) for transaction processing and multiple larger finalizer committees (FCs) for transaction finalization. To avoid system stagnation and thus guarantee stable performance, we ensure PSs’ liveness even if they are corrupted through the cooperation of PSs and FCs, thus eliminating the recovery process of corrupted PSs. To better trade-off security and scalability, we fine-tune the FCs to enable multiple FCs to coexist securely. As a result, DL-Chain allows a larger fraction of malicious nodes in each PS ($<1/2$) and thus can securely configure smaller shards for boosted stable and scalable concurrency. Evaluation results show that DL-Chain achieves up to 10 times improvement in throughput compared to existing solutions and provides stable concurrency with up to 2,550 nodes.
@@
Blockchain, blockchain sharding, concurrency
1 Introduction
@@
Blockchain sharding has attracted widespread attention as a technique to address low scalability in traditional blockchain [1, 2]. Its main idea is to partition the blockchain network into smaller groups, known as shards [21, 36, 19, 34, 23, 37, 20, 16]. Each shard manages a unique subset of the blockchain ledger state and performs intra-shard consensus to produce blocks concurrently. Generally, configuring a larger number of smaller shards tends to result in better transaction concurrency under the same network size [17, 31, 24].
@@
However, existing permissionless sharding systems require large shard sizes to ensure security, significantly limiting transaction concurrency in large-scale blockchain sharding systems [21, 6, 7]. In most permissionless blockchain sharding systems, nodes are randomly assigned to disjoint shards [21, 36, 20, 19, 23, 13]. This randomness causes smaller shards more likely to contain a larger fraction of malicious nodes that exceed the fault tolerance threshold (e.g., $\geq 1/3$ for BFT-typed intra-shard consensus mechanism), resulting in shard corruption [22] and compromised system security. Consequently, most current systems tend to configure large shard sizes (e.g., 600 nodes per shard in OmniLedger [21]) to substantially limit the probability of each shard’s corruption [21, 7, 6]. Unfortunately, such large shard sizes not only slow down intra-shard consensus but also decrease the network’s overall shard count, leading to reduced system transaction concurrency.
@@
While some previous studies have attempted to reduce shard size to improve concurrency, their solutions have various limitations. For instance, some works make less practical assumptions that the network is synchronous  [36, 20, 37]. Some works [16] need specific hardware [28], preventing widespread adoption. Some recent works [32, 17, 24] reduce shard sizes by allowing corrupted shards (i.e., shards with a larger fraction of malicious nodes). However, some of these works [32, 17] rely on network-wide consensus to ensure security, leading to limited scalability. Most importantly, those works [17, 24] frequently migrate the blockchain ledger to reshuffle the corrupted shards. Unfortunately, systems lose liveness during reshuffling, leading to severe temporary stagnation issues in real-world scenarios. For example, our measurements in Section 2.2 show that, when there are 24 shards, it takes over 100 minutes for a node to migrate pruned Ethereum [4] historical states.
@@
Figure 1: Illustration of Dual-Layer Architecture. A node simultaneously belongs to a PS and a FC.
This paper proposes DL-Chain, a highly concurrent blockchain sharding system with scalable and stable performance, to fill the aforementioned gap. To achieve the objectives, our core idea is a Dual-Layer architecture, as shown in Figure 1. In the lower layer, numerous smaller but more vulnerable proposer shards (PS) process transactions to achieve high concurrency. In the upper layer, secure finalizer committees (FC) safeguard potentially corrupted PSs to guarantee security. To guarantee scalability, we design multiple FCs in the upper layer, each responsible for safeguarding several disjoint PSs, rather than relying on network-wide consensus. More importantly, to ensure liveness and avoid stagnation, we design a cross-layer view change mechanism and fine-tune the quorum size within PSs, thus providing stable performance.
@@
Challenge 1: Balancing Security and Scalability with Dual-Layer Consensus. The most important challenge is how to ensure security despite the presence of corrupted shards while preserving scalability. To tackle this challenge, we propose a scalable Dual-Layer Consensus: Numerous smaller PSs process transactions and propose blocks through intra-shard consensus. However, their small size makes them more prone to corruption. Those corrupted PSs may fail to reach consensus (e.g., create forks) and jeopardize system security. Therefore, multiple larger FCs are set to verify and perform BFT-typed consensus on the consensus results of the corresponding PSs to finalize their transactions and eliminate forks. To balance scalability and security, we prudently fine-tune FC sizes to the minimum with negligible failure probability. This implies multiple FCs co-exist, and the proportion of malicious nodes in each FC is less than $1/3$ (i.e., this FC is secure and honest) with a high probability. As a result, each honest FC can provide finality for multiple PSs, mitigating network-wide consensus and enhancing scalability.
@@
However, it is difficult to achieve satisfactory performance and efficiency if the system only relies on the FCs in the Dual-Layer Consensus to ensure security (i.e., both liveness and safety). This is where the limitations of some existing layered architectures come into play [9, 13, 19, 11]. To address this challenge, we differentiate ourselves by requiring FCs and PSs to collaborate to guarantee liveness and safety, thus providing stable performance and reducing overhead. Detailed explanations are as follows.
@@
Challenge 2: Ensuring PSs’ Liveness for Stable Performance. The second challenge is how to maintain stable performance. In simplistic approaches [17, 24], corrupted shards that lose liveness require replacement or reshuffling. This results in time-consuming cross-shard state migration [13] and significant system stagnation issues. To tackle this challenge, we ensure that every PS (even when it is corrupted) will not lose liveness with FC’s help. We identify two scenarios wherein a corrupted shard may lose liveness. Firstly, a malicious leader may cause a PS to lose liveness by not proposing blocks. In such instances, a corrupted PS cannot rely on itself to perform the view change [12] to replace the malicious leader. To address this, we propose a cross-layer view change protocol that uses FCs to replace the malicious leaders for PSs. Secondly, when the number of honest nodes in a shard falls below the quorum size (the number of votes required for reaching consensus), malicious nodes can cause the shard to lose liveness by remaining silent. Consequently, we configure
#$\#\ of\ honest\ nodes\geq quorum\ size$ (1)
@@
Challenge 3: Ensuring PS’s Safety with Low Overhead. The third challenge is how to achieve low overhead for cross-layer communication in Dual-Layer Consensus while ensuring safety. When we rely exclusively on FCs for safety, FCs must obtain entire blocks from PSs to verify raw transactions and solve the forking issue, leading to considerable overhead. To address this challenge, our key idea is to rely on the consensus within PSs to guarantee the validity of raw transactions and utilize FCs to resolve the forking issue of PSs, so that only headers are transmitted across layers, thus reducing overhead. We observe that if the number of malicious nodes is less than the quorum size, a block containing invalid transactions will not pass the intra-shard consensus, as it cannot collect enough votes (honest nodes do not vote for invalid blocks). Therefore, our solution is to guarantee
$quorum\ size>\#\ of\ malicious\ nodes$ (2)
in each PS (guarantee transaction validity), guided by rigorous theoretical calculations. However, a corrupted PS can still fork its chain in this case. Luckily, such a safety attack can be detected by checking block headers. Therefore, we require FCs to obtain block headers from the corresponding PSs, select one fork branch for each PS, and finalize it through consensus. These designs ensure safety with low overhead.
@@
  • We propose DL-Chain, a high concurrency block-chain sharding system with scalable and stable performance. DL-Chain does not need network-wide consensus and prevents temporary stagnation.
  • To balance security and scalability, we propose scalable Dual-Layer Consensus. Each PS guarantees its liveness with FC’s help to provide stable performance. To reduce overhead, PSs and FCs synergize to guarantee safety so that only headers are transmitted across layers.
  • DL-Chain allows a larger fraction of malicious nodes in each PS. By combining the two aforementioned expressions 1 and  2, we derive the optimal tolerance for the fraction of malicious nodes in each PS ($<1/2$, instead of $<1/3$). This allows DL-Chain to configure smaller PSs, enhancing concurrency.
  • We implement DL-Chain based on Harmony [6], a well-known blockchain sharding project once had a top 50 market cap in cryptocurrency. Experimental results with up to 2,550 nodes show that DL-Chain can improve throughput 10x and maintain stable performance when the baseline system GearBox (CCS 22) [17] stagnates.
2 Background and Related Work
2.1 Background on Blockchain Sharding
Sharding has been extensively studied as a promising solution for improving the scalability of blockchain [26, 21, 36, 8, 16, 34, 23]. Its core idea is partitioning nodes into groups (aka shards). Each shard maintains a disjoint subset of the states and reaches intra-shard consensus to process disjoint transactions in parallel. A blockchain sharding system usually has the following main components.
@@ 2.1 Background on Blockchain Sharding
2) Intra-shard Consensus: After shard formation, each shard makes intra-shard consensus to append blocks into its shard chain. Most systems [36, 19, 23, 17] adopt BFT-typed consensus protocol (e.g., Practical Byzantine Fault Tolerance, PBFT) to produce blocks [33]. In such consensus protocols, honest nodes vote for valid blocks and will only accept blocks for which quorum size of votes have been collected.
@@
DL-Chain applies BFT-typed intra-shard consensus protocol, as many other sharding systems do. The BFT-typed intra-shard consensus protocol is a pluggable component. For implementation simplicity and fair comparison, we use the Fast Byzantine Fault Tolerance (FBFT) consensus protocol proposed by Harmony [6] in this paper. FBFT is a variation of PBFT [12], a leader-based consensus protocol providing the same security guarantee as PBFT. It, by default, withstands $<m/3$ malicious nodes with $12m/3+$ 1 quorum size in a group of size $m$ under a partial-synchronized network.
@@
3) Cross-shard Mechanism: Sharding partitions the ledger among shards, necessitating cross-shard transaction mechanisms to update each shard atomically. DL-Chain applies the relay-based mechanism similar to Monoxide [34] and Harmony [6], which initially packages cross-shard transactions in the source shard to deduct and forwards them to the destination shard with deduction proof. Then, the destination shard does the deposit operations. This protocol provides eventual atomicity and asynchronously lock-free processing to cross-shard transactions, preserving scalability even if almost all transactions are cross-shards [34].
2.2 Blockchain Sharding for Improved Concurrency
While sharding can enhance transaction concurrency through configuring more shards with small sizes, traditional systems [26, 21] favor larger shards for the security of each shard at the cost of concurrency. This is due to their reliance on each shard’s honesty, which is more probable with a larger size. While some works [36, 16, 19, 20, 37] aim to enhance concurrency by decreasing shard sizes, they have limitations. Some works [36, 20, 37] presume a synchronous network within each shard, impractical in large-scale blockchains. Others [16] leverage trusted hardware, requiring additional overhead per node. Lastly, some works [19, 23] achieve smaller shard sizes at the cost of overall system resilience.
Only a few existing studies improve concurrency by permitting corrupted shards. Free2Shard [32] leverages network-wide consensus and dynamic node allocation algorithm for security. However, it has an impractical assumption of a global synchronous network and relies on PoW consensus, which lacks deterministic finality. GearBox [17] ensures safety at the cost of liveness. It leverages a network-wide consensus to monitor each shard’s liveness and reshuffles unresponsive shards frequently. However, the dependence on consensus across the entire network impedes the system’s scalability. When reshuffling, the system discards corrupted shards, and services are temporarily unavailable until the state is migrated to an uncorrupted shard. CoChain [24] involves monitoring each shard by a group of other shards and can securely replace a corrupted shard when the number of corrupted shards is $F$ and each group has more than $13F+$ 1 shards. This implies that the system cannot tolerate any corrupted shards when the number of shards is no greater than 3. Moreover, like GearBox, CoChain faces temporary stagnation due to migrating historical ledgers and transactions from corrupted shards for recovery and processing.

### v2

# StableShard: Stable and Scalable Blockchain Sharding with High Concurrency via Collaborative Committees

@@ Abstract
Sharding enhances blockchain scalability by partitioning nodes into multiple groups for concurrent transaction processing. Configuring a large number of small shards usually helps improve transaction concurrency, but it also increases the fraction of malicious nodes in each shard, easily causing shard corruption and jeopardizing system security. Existing works attempt to improve concurrency by reducing shard sizes while maintaining security, but typically rely on time-consuming recovery of corrupted shards to restore liveness and network-wide consensus. This causes severe system stagnation and limits scalability.
To address this, we present StableShard, a sharded blockchain that securely provides high concurrency with stable and scalable performance. The core idea is to carefully co-design the division of labor between proposer shards (PSs) and finalizer committees (FCs): we deliberately assign 1) asymmetric roles and 2) matching parameters to PSs and FCs. Small PSs focus on fast transaction proposal and local validity, while large FCs focus on resolving forks, finalizing PS blocks, and maintaining liveness for faulty PSs via a cross-layer view-change protocol. Moreover, by fine-tuning key system parameters (e.g., shard size, quorum size), we ensure each PS to tolerate <1/2 fraction of malicious nodes without lossing liveness, and allow multiple FCs to securely coexist (each with <1/3 fraction of malicious nodes) for better scalability. Consequently, StableShard can safely configure many smaller PSs to boost concurrency, while FCs and PSs jointly guarantee safety and liveness without system stagnation, leading to stable and scalable performance. Evaluations show that StableShard achieves up to $10{\times}$ higher throughput than existing solutions and significantly more stable concurrency under attacks.
@@
Blockchain sharding, concurrency, in-place liveness, stable performance, division-of-labor and collaboration.
I Introduction
@@
Blockchain sharding has attracted widespread attention as a technique to address low scalability in traditional blockchain [bitcoin, eth]. Its main idea is to partition the blockchain network into smaller groups, known as shards [hellings2021byshard, cheng2024shardag, tao2020sharding, tao2023sharding, zheng2021meepo, amiri2021sharper, ruan2021blockchains, omniledger, rapidchain, pyramid, monoxide, li2022jenga, cycledger, repchain, sgxSharding]. Each shard manages a unique subset of the blockchain ledger state and performs intra-shard consensus to produce blocks concurrently. Generally, configuring a larger number of smaller shards tends to result in better transaction concurrency under the same network size [gearbox, instachain, cochain].
@@
However, existing permissionless sharding systems require large shard sizes to ensure security, significantly limiting transaction concurrency in large-scale blockchain sharding systems [omniledger, Harmony, zilliqa]. In most permissionless blockchain sharding systems, nodes are randomly assigned to disjoint shards [omniledger, rapidchain, repchain, pyramid, li2022jenga, sschain]. This randomness causes smaller shards more likely to contain a larger fraction of malicious nodes that exceed the fault tolerance threshold (e.g., $\geq 1/3$ for BFT-typed intra-shard consensus mechanism), resulting in shard corruption [beyondOneThird] and compromised system security. Consequently, current systems tend to configure large shard sizes (e.g., 600 nodes per shard in OmniLedger [omniledger]) to substantially limit the probability of each shard’s corruption [omniledger, zilliqa, Harmony]. Unfortunately, such large shard sizes not only slow down intra-shard consensus but also decrease the network’s overall shard count, leading to reduced system transaction concurrency.
@@
While some previous studies have attempted to reduce shard size to improve concurrency, their solutions have various limitations. Some works make less practical assumptions (e.g., synchronous network [rapidchain, repchain, cycledger, li2020polyshard, li2025sp], reduced resiliency [pyramid, li2022jenga], small scale [amiri2021sharper, hellings2021byshard, das2020efficient, zheng2021meepo, tao2023sharding, tao2020sharding, cheng2024shardag]) or require specialized hardware [sgxSharding, SGX], limiting practicality and deployability. More closely related, recent designs reduce shard sizes by tolerating corrupted shards (i.e., shards with a larger fraction of malicious nodes). However, they typically recover by triggering explicit recovery events after liveness breaks (e.g., replacing a faulty shard [cochain] or resampling a shard with larger parameters after detecting liveness failure [gearbox]). Unfortunately, the systems lose liveness during recovery, leading to severe temporary stagnation issues in real-world scenarios, and they are fundamentally different from keeping every small shard continuously live in-place. For example, our measurements in Section II-B show that with 24 shards, migrating pruned Ethereum [chainDataETH] historical states can take over 100 minutes per node. Moreover, some of these works [rana2022free2shard, gearbox] rely on network-wide consensus to ensure security, limiting scalability.
@@
Figure 1: Illustration of the Division-of-Labor Collaboration. A node simultaneously belongs to a PS and a FC.
This paper proposes StableShard, a highly concurrent blockchain sharding system that fills the above gap: it provides stable and scalable performance under corrupted shards without network-wide consensus and without any recovery procedure that leads to system stagnation. Our core idea is a Division-of-Labor Collaboration (DoLC) protocol, as shown in Figure 1. We co-design how proposer shards (PSs) and finalizer committees (FCs) divide responsibilities and collaborate so that safety and liveness are guaranteed by complementary mechanisms. Concretely, we deliberately assign asymmetric roles and matching parameters to PSs and FCs. Many small PSs are optimized for concurrency: they propose blocks quickly and guarantee transaction validity through a carefully chosen intra-shard quorum. Multiple larger FCs are optimized for robustness and scalability: each FC finalizes multiple disjoint PSs by running lightweight BFT on headers to resolve forks and provide finality, avoiding network-wide consensus. Crucially, FCs also maintain stagnation-free liveness for faulty PSs via a cross-layer view-change protocol, so that even corrupted PSs remain live in-place instead of triggering recovery events. Moreover, by fine-tuning key parameters (e.g., committee sizes, quorum thresholds) jointly, StableShard enables PSs to tolerate a Byzantine fraction approaching <1/2<1/2 while keeping each FC secure (<1/3<1/3) with overwhelmingly high probability, thereby supporting many small PSs and multiple coexisting FCs for stable and scalable concurrency.
Realizing the above goal requires addressing the following challenges.
@@
Challenge 1: Scalable Finalization without Network-wide Consensus. The first challenge is how to ensure security despite the presence of corrupted shards while preserving scalability. To tackle this challenge, we propose a scalable PS-FC Division-of-Labor Collaboration (DoLC) protocol: Numerous smaller PSs process transactions and propose blocks through intra-shard consensus. However, their small size makes them more prone to corruption. Those corrupted PSs may fail to reach consensus (e.g., create forks) and jeopardize system security. Therefore, multiple larger FCs are set to verify and perform BFT-typed consensus on the consensus results of the corresponding PSs to finalize their transactions and eliminate forks (via our fork selection rules). To balance scalability and security, we prudently fine-tune FC sizes to the minimum with negligible failure probability. This implies multiple FCs co-exist, and the proportion of malicious nodes in each FC is less than $1/3$ (i.e., this FC is secure and honest) with a extremely high probability. As a result, each honest FC can provide finality for multiple PSs, mitigating network-wide consensus and enhancing scalability.
@@
However, it is difficult to achieve stable performance and efficiency if the system only relies on the FCs to ensure security (i.e., both liveness and safety). This is where the limitations of some existing works come into play [sschain, pyramid, gearbox, cochain, rana2022free2shard]. To address this challenge, we differentiate ourselves by requiring FCs and PSs to divide responsibilities and collaborate to guarantee liveness and safety, thus providing stable performance and reducing overhead. Detailed explanations are as follows.
@@
Challenge 2: Stagnation-free Liveness without Recovery for Stable Performance. The second challenge is how to maintain stable performance. In simplistic approaches [gearbox, cochain], corrupted shards that lose liveness require replacement or reshuffling. This results in time-consuming cross-shard state migration [sschain] and significant system stagnation issues. To tackle this challenge, we ensure that every PS (even when it is corrupted) will not lose liveness with FC’s help and quorum co-design. We identify two scenarios wherein a corrupted shard may lose liveness. Firstly, a malicious leader may cause a PS to lose liveness by not proposing blocks. In such instances, a corrupted PS cannot rely on itself to perform the view change [pbft] to replace the malicious leader. To address this, we propose a cross-layer view change protocol that uses FCs to replace the malicious leaders for PSs. Secondly, when the number of honest nodes in a shard falls below the quorum size (the number of votes required for reaching consensus), malicious nodes can cause the shard to lose liveness by remaining silent. Consequently, we deliberately configure
$of\ honest\ nodes\geq quorum\ size$ (1)
@@
Challenge 3: Low-overhead Safety via Complementary Responsibilities. The third challenge is how to achieve low overhead for cross-layer communication in DoLC while ensuring safety. When we rely exclusively on FCs for safety, FCs must obtain entire blocks from PSs to verify raw transactions and solve the forking issue, leading to considerable overhead. To address this challenge, our key idea is to rely on the consensus within PSs to guarantee the validity of raw transactions and utilize FCs to resolve the forking issue of PSs, so that only headers are transmitted across layers, reducing overhead. We observe that if the number of malicious nodes is less than the quorum size, a block containing invalid transactions will not pass the intra-shard consensus, as it cannot collect enough votes (honest nodes do not vote for invalid blocks). Therefore, our solution is to guarantee
quorumsize>#ofmaliciousnodesquorum\ size>\#\ of\ malicious\ nodes (2)
in each PS (guarantee transaction validity), guided by careful parameterization and rigorous theoretical calculations. However, a corrupted PS can still fork its chain in this case. Luckily, such a safety attack can be detected by checking block headers. Therefore, we require FCs to obtain block headers from the corresponding PSs, select one fork branch for each PS, and finalize it through consensus. These designs ensure safety with low overhead.
@@
  • We present StableShard, a sharded blockchain that achieves high concurrency with stable and scalable performance. StableShard eliminates recovery procedures that cause stagnation and avoids network-wide consensus.
  • We co-design a scalable PS-FC Division-of-Labor Collaboration (DoLC) protocol. Small PSs maximize concurrency and guarantee transaction validity via quorum voting. To provide stable performance, FCs maintain in-place liveness for faulty PSs via a cross-layer view-change mechanism. For balanced security and scalability, multiple large and secure FCs provide header-only fork resolution and finality for multiple PSs.
  • We introduce a joint parameterization (committee sizes and quorum thresholds). By jointly satisfying expressions 1 and  2, and together with the DoLC protocol, StableShard tolerates <1/2<1/2 (instead of <1/3<1/3) fraction of malicious nodes per PS while guaranteeing stagnation-free liveness. It also enables multiple secure FCs (<1/3<1/3 fraction of malicious nodes) to coexist. Jointly, StableShard achieves high concurrency, stable and scalable performance.
  • We implement StableShard based on Harmony [Harmony], an industrial sharding codebase once had a top 50 market cap in cryptocurrency, and evaluate it at scale (up to 2,550 nodes). Results show up to $10\times$ throughput improvement and significantly more stable concurrency under attacks over the baseline system GearBox (CCS 22) [gearbox].
II Background and Related Work
II-A Background on Blockchain Sharding
Sharding has been extensively studied as a promising solution for improving the scalability of blockchain [hellings2021byshard, cheng2024shardag, tao2020sharding, tao2023sharding, zheng2021meepo, amiri2021sharper, ruan2021blockchains, omniledger, rapidchain, pyramid, monoxide, li2022jenga, cycledger, repchain, sgxSharding]. Its core idea is partitioning nodes into groups (aka shards). Each shard maintains a disjoint subset of the states and reaches intra-shard consensus to process disjoint transactions in parallel. A blockchain sharding system usually has the following main components.
@@ II-A Background on Blockchain Sharding
2) Intra-shard Consensus: After shard formation, each shard makes intra-shard consensus to append blocks into its shard chain. Most systems [cheng2024shardag, zheng2021meepo, rapidchain, pyramid, li2022jenga, gearbox] adopt BFT-typed consensus protocol (e.g., Practical Byzantine Fault Tolerance, PBFT) to produce blocks [sok]. In such consensus protocols, honest nodes vote for valid blocks and will only accept blocks for which quorum size of votes have been collected.
@@
StableShard applies BFT-typed intra-shard consensus protocol, as many other sharding systems do. The BFT-typed intra-shard consensus protocol is a pluggable component. For implementation simplicity and fair comparison, we use the Fast Byzantine Fault Tolerance (FBFT) consensus protocol proposed by Harmony [Harmony] in this paper. FBFT is a variation of PBFT [pbft], a leader-based consensus protocol providing the same security guarantee as PBFT. It, by default, withstands <m/3<m/3 malicious nodes with $2m/3+1$ quorum size in a group of size $m$ under a partial-synchronized network. Moreover, like existing BFT-type consensus protocols in partially synchronous networks, FBFT also incorporates a timeout and view change mechanism to ensure consensus liveness. Similarly, during the asynchronous phase before GST, the timeout duration gradually increases with each view change and eventually converges to a stable value (once the network becomes synchronous), to accommodate the asynchronous nature of the network and prevent overly frequent view changes [pbft, hotstuff, pyramid].
@@
3) Cross-shard Mechanism: Sharding partitions the ledger among shards, necessitating cross-shard transaction mechanisms to update each shard atomically. StableShard applies the relay-based mechanism similar to Monoxide [monoxide] and Harmony [Harmony], which initially packages cross-shard transactions in the source shard to deduct and forwards them to the destination shard with deduction proof. Then, the destination shard does the deposit operations. This protocol provides eventual atomicity and asynchronously lock-free processing to cross-shard transactions, preserving scalability even if almost all transactions are cross-shards [monoxide].
II-B Blockchain Sharding for Improved Concurrency
We group the most relevant prior work into three categories according to how they trade concurrency, safety, and liveness, and how they handle corrupted shards and performance stability.
1) Permissionless sharding with conservative (large) shards. Classic permissionless sharding systems (e.g., [elastico, omniledger, Harmony]) improve concurrency by parallelizing intra-shard consensus, but typically require large shard sizes so that each shard remains honest with overwhelming probability. This conservative choice preserves safety/liveness under the standard <1/3<1/3 BFT threshold, yet it limits scalability because large shards slow consensus and reduce the number of shards, thus capping concurrency.
2) Small-shard sharding under stronger assumptions or weakened resilience. A separate line of work reduces shard sizes via additional assumptions or auxiliary support, e.g., assuming synchrony within shards [rapidchain, repchain, cycledger, li2020polyshard, li2025sp] or leveraging trusted hardware [sgxSharding]. Some works [pyramid, li2022jenga] achieve smaller shard sizes at the cost of overall system resilience. Some other designs achieve very small shards by restricting the overall network scale or not rigorously accounting for corruption probability at large scale [amiri2021sharper, hellings2021byshard, das2020efficient, zheng2021meepo, tao2023sharding, tao2020sharding, cheng2024shardag]. These approaches can boost concurrency, but their assumptions (synchrony, extra hardware, weakened resilience, or small deployment scale) reduce practicality in large permissionless settings.
3) Corrupted-shard-tolerant designs and hierarchical/decoupled confirmations. Only a few blockchain sharding systems explicitly permit corrupted shards to shrink committees while attempting to preserve overall security. Free2Shard [rana2022free2shard] relies on network-wide coordination (and PoW-style assumptions), which limits deterministic finality and can constrain scalability. GearBox [gearbox] leverages the safety-liveness dichotomy and detects liveness violations to trigger "gear switching" (i.e., adjusting committee parameters, injecting more nodes) so that liveness can be restored. However, such recovery is orchestrated at a network-wide level and can introduce instability under repeated or adaptive attacks, as shards may frequently enter the recovery path. Similarly, systems that rely on network-wide coordination [sschain, rana2022free2shard] tend to face scalability limits due to global monitoring/consensus overhead. CoChain [cochain] monitors each shard using groups of other shards and recovers the system via shard replacement/takeover once a corrupted shard is found (Consensus-on-Consensus). While effective, this recovery model is fundamentally different from keeping small shards continuously live in-place, and it still raises practical costs for transferring historical ledgers/transaction context to the replacement executor.
Severe Stagnation from Recovery/State Movement. A common Achilles’ heel of many corrupted-shard-tolerant designs is that liveness recovery is coupled with heavy recovery actions (e.g., state/ledger transfer, reshuffling, executor replacement). Even when safety can be maintained, these recovery paths may pause service and cause large latency spikes, which is unacceptable for stable performance.